# Event-based Sensor Fusion and Application on Odometry: A Survey

### v2

@@ III-E Fusion with LiDARs
Some of the research is still in the early stages, with a primary focus on achieving automatic and accurate calibration. In [47], the authors introduce a direct extrinsic calibration method between event cameras and LiDAR, exploiting the high temporal resolution and dynamic range of event cameras. This technique eliminates the need for frame-based intermediaries and optimizes 6-DoF extrinsic calibration through information-based correlation methods, improving sensor alignment. Similarly, in [48], a novel calibration method is proposed, using a 3D marker to generate stable patterns detectable in both 2D point sets and 3D point clouds. This approach facilitates accurate calibration between the event camera and LiDAR by leveraging light changes to trigger event detection, which is crucial for achieving stable fusion. This paves the way for the fusion of event cameras and LiDARs.
Beyond calibration, LiDAR and event camera fusion can enhance depth perception. In [49], a method proposes fusing sparse LiDAR and event data to generate dense depth maps, improving spatial perception. This improved point cloud map can further facilitate the LiDAR-based odometry. In the work [50], the authors discovered that data from event cameras share a homogeneous nature with RGB and LiDAR data in both visual and motion spaces. This characteristic makes event cameras particularly suitable for bridging the gap between RGB and LiDAR in motion fusion or scene flow tasks. The finding highlights the potential for a hierarchical fusion framework that effectively integrates these sensors, making the fusion of event cameras with other sensors like LiDAR and conventional cameras highly promising.
IV Event Camera Dataset
Unlike other reviews of event camera-based datasets, this part of the paper aims to highlight several event camera-related datasets that have particularly been developed to support research in event-based sensor fusion for odometry, enabling accurate pose estimation and motion tracking in dynamic environments. The datasets, along with the sensors involved, are listed in Table II.
@@ V Conclusion and Discussion
V-A Conclusion
Due to their high temporal resolution, low latency, high dynamic range, and reduced motion blur compared to traditional cameras, event cameras offer significant potential benefits for odometry calculation by capturing precise changes in brightness. Unlike previous review papers focused on event cameras, this paper provides a comprehensive and detailed analysis of the application of event cameras for odometry, exploring various sensor fusion approaches. These include event camera-only systems, fusion with IMUs, stereo event camera setups, integration with traditional cameras, and combinations with LiDAR. Additionally, we briefly discuss event camera-related datasets that are relevant for advancing sensor fusion, particularly in odometry applications.
V-B Discussion
Event camera holds significant potential for future applications across various domains. Its role in odometry or pose estimation has been specifically addressed in the paper. In Visual Odometry (VO) or Visual-Inertial Odometry (VIO), the use of event cameras can effectively mitigate motion blur and reduce the challenges posed by low-light environments, thanks to their wider dynamic range. For LiDAR-based odometry, similar to their application with conventional cameras, event cameras can help reduce drift by compensating for the loss of geometric information in LiDAR point clouds with photometric information. Additionally, event cameras can be integrated with LiDAR-generated images, as modern LiDAR systems often produce dense, low-resolution panoramic images with encoded infrared light reflectivity and intensity shown in the bottom of Figure 1. This integration could further enhance the robustness and accuracy of odometry in complex environments. Although the current stage of fusion with LiDAR is still relatively early, it shows promising potential for future developments.
Acknowledgment
This research is supported by the Research Council of Finland’s Digital Waters (DIWA) flagship (Grant No. 359247) as well as the DIWA Doctoral Training Pilot project funded by the Ministry of Education and Culture (Finland).

### v3

@@ IV-E Fusion with LiDARs
Some of the research is still in the early stages, with a primary focus on achieving automatic and accurate calibration. In [48], Ta et al. introduced a direct extrinsic calibration method between event cameras and LiDAR, exploiting the high temporal resolution and dynamic range of event cameras. This technique eliminates the need for frame-based intermediaries and optimizes 6-DoF extrinsic calibration through information-based correlation methods, improving sensor alignment. Similarly, Song et al. [49] proposed a novel calibration method using a 3D marker to generate stable patterns detectable in both 2D point sets and 3D point clouds. This approach facilitates accurate calibration between the event camera and LiDAR by leveraging light changes to trigger event detection, which is crucial for achieving stable fusion. This paves the way for the fusion of event cameras and LiDARs.
Beyond calibration, LiDAR and event camera fusion can enhance depth perception. Cui et al. [50] proposed a method to fuse sparse LiDAR and event data to generate dense depth maps, improving spatial perception. This improved point cloud map can further facilitate the LiDAR-based odometry. In [51], Zhou et al. discovered that data from event cameras share a homogeneous nature with RGB and LiDAR data in both visual and motion spaces. This characteristic makes event cameras particularly suitable for bridging the gap between RGB and LiDAR in motion fusion or scene flow tasks. The finding highlights the potential for a hierarchical fusion framework that effectively integrates these sensors, making the fusion of event cameras with other sensors like LiDAR and conventional cameras highly promising.
IV-F Event Camera Datasets
With the increasing interest towars event cameras and their utilisation in robotic odometry, many datasets has been introduced in recent years. In Table II,s we list those datasets that have particularly been developed to support research in event-based sensor fusion for odometry, enabling accurate pose estimation and motion tracking in dynamic environments.
@@ V Conclusion and Discussion
Due to their high temporal resolution, low latency, high dynamic range, and reduced motion blur compared to traditional cameras, event cameras offer significant potential benefits for odometry calculation by capturing precise changes in brightness. Unlike previous review papers focused on event cameras, this survey provides a comprehensive and detailed analysis of the application of event cameras for odometry, exploring various sensor fusion approaches. These include event camera-only systems, fusion with IMUs, stereo event camera setups, integration with traditional cameras, and combinations with LiDAR. Additionally, we briefly discuss event camera-related datasets that are relevant for advancing sensor fusion, particularly in odometry applications.
Event camera holds significant potential for future applications across various domains. Its role in odometry or pose estimation has been specifically addressed in the paper. In Visual Odometry (VO) or Visual-Inertial Odometry (VIO), the use of event cameras can effectively mitigate motion blur and reduce the challenges posed by low-light environments, thanks to their wider dynamic range. For LiDAR-based odometry, similar to their application with conventional cameras, event cameras can help reduce drift by compensating for the loss of geometric information in LiDAR point clouds with photometric information. Additionally, event cameras can be integrated with LiDAR-generated images, as modern LiDAR systems often produce dense, low-resolution panoramic images with encoded infrared light reflectivity and intensity as shown in the bottom of Fig. 1. This integration could further enhance the robustness and accuracy of odometry in complex environments. Although the current stage of fusion with LiDAR is still relatively early, it shows promising potential for future developments.